If the product of two consecutive even numbers is increased by 28, it will equal the square of the larger number. Find the two numbers.
Answer by mananth(16946)

let the numbers be x & x+2
x(x+2)+28 =(x+2)^2
x^2+2x+28=x^2+4x+4
2x=24
x=12
12 & 14 are the numbers